# Advancing Semantic Future Prediction through Multimodal Visual Sequence Transformers

### v1

@@ Appendix B Additional Qualitative Results
In Figs. 7, 8, 9, 10 and 11, we present additional qualitative results using FUTURIST for forecasting semantic segmentation and depth maps over extended time horizons. These results were generated through autoregressive rollouts. Starting with a sequence of four context frames ($X_{t-9}$ to $X_{t}$), the model predicts up to 48 future frames, corresponding to 2.88 seconds, with a frame interval of 3.
The examples in Figs. 7, 8, 9 and 10 demonstrate that the model effectively preserves temporal coherence, maintaining consistent relationships between static and dynamic elements over time. The depth maps transition smoothly between frames, aligning well with changes in the scene. For instance, Fig. 7 and Fig. 9 show complex scenes with numerous static and moving objects. Here, the model captures the ego vehicle’s motion accurately, enabling precise predictions. In Fig. 8, the model successfully predicts the motion of a car crossing perpendicularly to the ego vehicle, while Fig. 10 highlights the model’s ability to anticipate the completion of a right turn.
However, as shown in Fig. 11, the quality of predictions degrades toward the end of the rollout, particularly in the final four frames. The car masks become elongated instead of approaching the ego vehicle as expected. This degradation likely stems from a mismatch between training and inference: during training, the model uses teacher forcing with oracle-provided context frames, whereas at inference, it relies on its own noisier predictions from previous steps. As discussed in Appendix C, future work should aim to address this issue.
@@ Appendix C Limitations and Future Work
[…]

### v2

@@ Appendix B Additional Qualitative Results
In  Figures 7, 8, 9, 10 and 11, we present additional qualitative results using FUTURIST for forecasting semantic segmentation and depth maps over extended time horizons. These results were generated through autoregressive rollouts. Starting with a sequence of four context frames ($X_{t-9}$ to $X_{t}$), the model predicts up to 48 future frames, corresponding to 2.88 seconds, with a frame interval of 3.
The examples in  Figures 7, 8, 9, and 10 demonstrate that the model effectively preserves temporal coherence, maintaining consistent relationships between static and dynamic elements over time. The depth maps transition smoothly between frames, aligning well with changes in the scene. For instance, Figure 7 and Figure 9 show complex scenes with numerous static and moving objects. Here, the model captures the ego vehicle’s motion accurately, enabling precise predictions. In Figure 8, the model successfully predicts the motion of a car crossing perpendicularly to the ego vehicle, while Figure 10 highlights the model’s ability to anticipate the completion of a right turn.
However, as shown in Figure 11, the quality of predictions degrades toward the end of the rollout, particularly in the final four frames. The car masks become elongated instead of approaching the ego vehicle as expected. This degradation likely stems from a mismatch between training and inference: during training, the model uses teacher forcing with oracle-provided context frames, whereas at inference, it relies on its own noisier predictions from previous steps. As discussed in Appendix C, future work should aim to address this issue.
@@ Appendix C Limitations and Future Work
Another promising direction is extending FUTURIST to instance and panoptic segmentation tasks. Our approach could be adapted by encoding instance information through pixel-wise offsets to instance centers [51, 54, 52] rather than using instance IDs directly. This would allow each pixel within an instance mask to be represented by its x and y offsets to the corresponding instance center. These offsets could be processed as an additional modality with a dedicated embedding layer. During inference, instance masks could be recovered from these predicted offsets using a Hough transform-like approach and combined with semantic predictions to generate panoptic segmentation. This extension would enhance the applicability of our method to scenarios requiring instance-level understanding.
@@
Figure 9: Long-term semantic segmentation and depth predictions for Scene: Frankfurt $\textbf{(0\_[1217-1246])}.$ Our model accurately preserves the relationships between static and dynamic elements and the motion of the ego vehicle.